# Seeing like a Cephalopod: Colour Vision with a Monochrome Event Camera

### v1

@@ Abstract
Cephalopods exhibit unique colour discrimination capabilities despite having one type of photoreceptor, relying instead on chromatic aberration induced by their ocular optics and pupil shapes to perceive spectral information. We took inspiration from this biological mechanism to design a spectral imaging system that combines a ball lens with an event-based camera. Our approach relies on a motorised system that shifts the focal position, mirroring the adaptive lens motion in cephalopods. This approach has enabled us to achieve wavelength-dependent focusing across the visible light and near-infrared spectrum, making the event a spectral sensor. We characterise chromatic aberration effects, using both event-based and conventional frame-based sensors, validating the effectiveness of bio-inspired spectral discrimination both in simulation and in a real setup as well as assessing the spectral discrimination performance. Our proposed approach provides a robust spectral sensing capability without conventional colour filters or computational demosaicing. This approach opens new pathways toward new spectral sensing systems inspired by nature’s evolutionary solutions. Code and analysis are available at: https://samiarja.github.io/neuromorphic_octopus_eye/.
@@ 1 Introduction
Many shallow-water cephalopods have evolved off-axis or complex pupil shapes (e.g., U- or W-shaped pupils). The combination of these specialized pupils and ball lenses projects distinct spatial patterns onto the retina, depending on the spectral content and position of incoming light. Elongated or semi-annular pupils collect light at oblique angles, intentionally enhancing chromatic blur and spectral cues, though at the expense of spatial sharpness [27], prioritizing spectral discrimination over spatial resolution.
@@ 6 Conclusion
In this paper, we presented a cephalopod-inspired approach that leverages chromatic aberration from a ball lens to capture spectral information across the visible and near-infrared ranges. By mimicking cephalopods’ dynamic focal shifts through a motorized system, we demonstrated how event-based and frame-based cameras can directly capture spectral data, eliminating the need for colour filters or complex demosaicing, even for sparse, narrowband signals. Our ray simulator, which accounts for the ball lens’s graded-index properties and the curved retina structure, was used to characterise the role of chromatic aberration and confirmed its ability to enable spectral discrimination. Using a motorised optical platform, we characterised the system’s ability for spectral discrimination and were able to infer spectral information directly from event stream data. Notably, the system’s spectral resolution fits the requirements imposed by cephalopods’ habitat. The aquatic attenuation spectrum supports the notion of higher spectral resolution at shorter wavelengths, where colour cues may convey information regarding the distance and composition of the scene, while higher photon sensitivity in long wavelengths can help penetrate the lower-intensity signals at these bands. This work signals new directions for visual sensing systems driven by nature’s evolutionary solutions and a novel approach to enhancing neuromorphic vision systems.

### v2

@@ Abstract
Cephalopods exhibit unique colour discrimination capabilities despite having one type of photoreceptor, relying instead on chromatic aberration induced by their ocular optics and pupil shapes to perceive spectral information. We took inspiration from this biological mechanism to design a spectral imaging system that combines a ball lens with an event-based camera. Our approach relies on a motorised system that shifts the focal position, mirroring the adaptive lens motion in cephalopods. This approach has enabled us to achieve wavelength-dependent focusing across the visible light and near-infrared spectrum, making the event a spectral sensor. We characterise chromatic aberration effects, using both event-based and conventional frame-based sensors, validating the effectiveness of bio-inspired spectral discrimination both in simulation and in a real setup, as well as assessing the spectral discrimination performance. Our proposed approach provides a robust spectral sensing capability without conventional colour filters or computational demosaicing. This approach opens new pathways toward new spectral sensing systems inspired by nature’s evolutionary solutions. Code and analysis are available at: Project Page.
@@ 1 Introduction
Many shallow-water cephalopods have evolved off-axis or complex pupil shapes (e.g., U- or W-shaped pupils). The combination of these specialised pupils and ball lenses projects distinct spatial patterns onto the retina, depending on the spectral content and position of incoming light. Elongated or semi-annular pupils collect light at oblique angles, intentionally enhancing chromatic blur and spectral cues, though at the expense of spatial sharpness [27], prioritising spectral discrimination over spatial resolution.
@@ 6 Conclusion
In this paper, we presented a cephalopod-inspired approach that leverages chromatic aberration from a ball lens to capture spectral information across the visible and near-infrared ranges. By mimicking cephalopods’ dynamic focal shifts through a motorised system, we demonstrated how event-based and frame-based cameras can directly capture spectral data, eliminating the need for colour filters or complex demosaicing, even for sparse, narrowband signals. Our ray simulator, which accounts for the ball lens’s graded-index properties and the curved retina structure, was used to characterise the role of chromatic aberration and confirmed its ability to enable spectral discrimination. Using a motorised optical platform, we characterised the system’s ability for spectral discrimination and were able to infer spectral information directly from event stream data. Notably, the system’s spectral resolution fits the requirements imposed by cephalopods’ habitat. The aquatic attenuation spectrum supports the notion of higher spectral resolution at shorter wavelengths, where colour cues may convey information regarding the distance and composition of the scene, while higher photon sensitivity in long wavelengths can help penetrate the lower-intensity signals at these bands. This work signals new directions for visual sensing systems driven by nature’s evolutionary solutions and a novel approach to enhancing neuromorphic vision systems.